Calculate the mass in g of potassium chloride and volume in dm3 of oxygen gas produced at S.T.P when 100g potassium chlorate is decomposed
1
Expert's answer
2021-08-18T17:08:03-0400

2KClO3(s) → 2KCl(s) + 3O2(g)

According to the equation

n (KClO3) = n (KCl)

n (O2) = 3/2 x n (KClO3) = 3/2 x n (KCl)

n = m / M

M (KClO3) = 122.5 g/mol

n (KClO3) = 100 / 122.5 = 0.8 mol

n (KCl) = n (KClO3) = 0.8 mol

M (KCl) = 74.6 g/mol

m (KCl) = n x M = 0.8 x 74.6 = 60.9 g

n (O2) = 3/2 x 0.8 = 1.2 mol

n (O2) = V/22.4

V (O2) = n (O2) x 22.4 = 1.2 x 22.4 = 26.9 l